OpenStudy (anonymous):

2. The general equation of a plane is Ax + By + Cz = D, where A, B, C, and D are real numbers and A is nonnegative. Find the equation of the plane containing the points (3, 0, 0), (0, 8, 0), and (0, 0, 6). Show each step of your process. Then graph the plane.

OpenStudy (anonymous):

Work this problem as 3 equations and 3 unknowns: Rewrite the standard equation like this: A/Dx + B/Dy + C/Dz = 1 To clean this up I used new variables: A/D = m, B/D = n, and C/D = o So: mx + ny + oz = 1 Now put each point onto this equation: 3m + 0 + 0 = 1 0 + 8n + 0 =1 0 + 0 + 6o = 1 In this case you can solve each directly, otherwise you could use substitution. m = 1/3 n = 1/8 o = 1/6 Put these values in the cleaned up equation: 1/3 x + 1/8 y + 1/6 z = 1 This is in standard form, however you can multiply each term by 24 to remove fractions: 8x + 3y + 4z = 24.

OpenStudy (anonymous):

does that work

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

no problem